Suppose the data represent the inches of rainfall in April for a certain city over the course of 20 years.

Determine the quartiles.

0.39

1.72

3.58

5.33

0.61

1.93

3.93

5.44

1.06

2.36

4.34

5.77

1.35

2.73

4.49

6.04

1.52

3.14

4.96

6.31

What are the​ quartiles?

Upper Q 1

equalsnothing

​(Round to three decimal places as​ needed.)
Answer #1

Quartiles are the values that divide a list of numbers into quarters 25% , 50% , 75%.

Number Values ( X )
1 0.61
2 1.06
3 1.35
4 1.52
5 1.72
6 1.93
7 2.36
8 2.73
9 3.14
10 3.58
11 3.93
12 4.34
13 4.49
14 4.96
15 5.33
16 5.44
17 5.77
18 6.04
19 6.31

Q1 = (1/4 n + 1/4 )th value
Q1 = (1/4 * 19 + 1/4 )th value
Q1 = 1.72

Q2 = (1/2 n + 1/2 )th value
Q2 = (1/2 * 19 + 1/2 )th value
Q2 = 3.58

Q3 = (3/4 n + 3/4 )th value
Q3 = (3/4 * 19 + 3/4 )th value
Q3 = 5.33